OpenStudy (anonymous):

You have 26 cards, each labeled with a letter of the alphabet. What is the probability of drawing a card that contains one of the letters in the word 'MATH'? A) 1/676 B) 2/13 C) 2/26 D) 3/26

OpenStudy (mathstudent55):

How many total cards (outcomes)? How many desired outcomes? P(event) = (number of desired outcomes) / (number of total outcomes)

OpenStudy (anonymous):

I know It's B

OpenStudy (word2):

There's is a 1/26 probability of choosing the "A", 1/26 for choosing the "T", and 1/26 for choosing the "H". In this question, we're looking at the probability of drawing one of the four cards in a set of 26 (M, A, T, and H). In this case we have four options that we want out of 26. So, the probability here would be the same as adding together the chances of drawing each card together (since we're more likely to draw one of four cards than a single card). 1/26 + 1/26 + 1/26 + 1/26

OpenStudy (mathstudent55):

Since there are 4 desired outcomes, M, A, T, and H, and 26 total possible outcomes, the probability is: P(drawing M, A, t, or H) = 4/26 = 2/13